# When do Reforms meet fairness concerns in school admissions?

### v3

@@ 6. Conclusions
We now prove that $\beta\geq\alpha$. We prove the claim that there is no student in $I\setminus\{i,i_{1},\ldots,i_{n+1}\}$ who is not a blocking student of $\nu$ but a blocking student of $\mu$. Let $j\notin I\setminus\{i,i_{1},\ldots,i_{n+1}\}$ and note that $\nu(j)=\mu(j)$. If student $j$ is matched under $\nu$, then she is not a blocking student of $\nu$ and $\mu$. Suppose that $\nu(j)=\emptyset$, and for a contradiction, that she is not a blocking student of $\nu$ but a blocking student of $\mu$ under $(P,\succ,q)$. There is a school $s^{\prime}$ such that $s^{\prime}\mathrel{P_{j}}\mu(j)$ and either (i) school $s^{\prime}$ has an empty seat under $\mu$ or (ii) there is a student $j^{\prime}$ such that $\mu(j^{\prime})=s^{\prime}$ and $j\mathrel{\succ_{s^{\prime}}j^{\prime}}$. Consider (i). As above, school $s^{\prime}$ has an empty seat under $\nu$. Therefore, student $j$ is also a blocking student of $\nu$, contradicting our assumption. Consider (ii). Suppose that school $s^{\prime}$ did not reject any student in any step from the step at which student $i$ applies to school $s$ to the end. Then $\mu(j^{\prime})=\nu(j^{\prime})=s^{\prime}$ and student $j$ is also a blocking student of $\nu$. Suppose that school $s^{\prime}$ has rejected a student $j^{\prime\prime}$ such that $\nu(j^{\prime\prime})=s^{\prime}$. Since $\mu(j^{\prime})=s^{\prime}$, then $j^{\prime}\mathrel{\succ_{s^{\prime}}}j^{\prime\prime}$. Thus $j\mathrel{\succ_{s^{\prime}}}j^{\prime\prime}$ and since $\nu(j^{\prime\prime})=s^{\prime}$, student $j$ is a blocking student of $\nu$, contradicting again our assumption. Therefore there is no student who is not a blocking student of $\nu$ but a blocking student of $\mu$ under $(P,\succ,q)$. Thus $\alpha\geq\beta$.
@@
By this claim, $\alpha+y\geq\beta+x$. Therefore, $\nu$ has weakly more number of blocking students than $\mu$. Starting from $N=\emptyset$ and successively replacing students preferences in any order we conclude that $GS(P^{k-1},\succ,q)$ has weakly more number of blocking students than $GS(P^{k},\succ,q)$.
@@
Finally, we describe a problem where the outcome of $GS^{\ell}$ has more blocking students than the outcome of $GS^{k}$. Let $(P,\succ,q)$ be a problem where each school has one seat, each student has $k$ acceptable schools and such that students have a common ranking of schools. Then, $GS^{k}(P,\succ,q)=GS(P,\succ,q)$. Thus $GS^{k}(P,\succ,q)$ is stable at $(P,\succ,q)$. Let $s$ be the school that students have ranked at the $k$’th position starting from the top. Since there are more students than schools and $k>\ell$, at least one student is not matched under $GS^{\ell}(P,\succ,q)$ and no student is matched to school $s$ even though every student prefers it to being unmatched. Therefore, there are more blocking students for $GS^{\ell}(P,\succ,q)$ than $GS^{k}(P,\succ,q)$ under $(P,\succ,q)$. ∎
Proof of Theorem 2.
Let $(P,\succ,q)$ be a problem. By Lemma 1, the matchings $SD(P^{\ell}_{N},P_{-N},\succ,q)$ and $SD(P^{k}_{N},P_{-N},\succ,q)$ are the unique Nash equilibrium outcomes of the games $(SD^{\ell},P)$ and $(SD^{k},P)$, respectively. Following the proof of Theorem 1, we show that the matching $SD(P^{\ell}_{N},P_{-N},\succ,q)$ has at least as many blocking students as the matching $SD(P^{k}_{N},P_{-N},\succ,q)$. ∎
@@
Part (i): Let $i$ be a blocking student of $\mu=\beta(P^{k},\succ,q)$. There is a school $s$ such that the pair $(i,s)$ blocks $\mu$ under $(P,\succ,q)$. Then, $s\mathrel{P_{i}}\mu(i)$ and either (a) school $s$ has an empty seat under $\mu$ or (b) there is a student $j$ such that $\mu(j)=s$ and $i\mathrel{\succ_{s}}j$. We claim that student $i$ did not rank school $s$ first under $P_{i}$. Otherwise, school $s$ has rejected student $i$ at the first step of the Boston algorithm under $(P^{k},\succ,q)$. This is because $k>1$ and the top ranked schools are considered under $\beta^{k}$. This contradicts the assumption that school $s$ has an empty seat or has accepted student $j$ with $i\mathrel{\succ_{s}}j$. Let $P^{s}_{i}$ be a preference relation where $i$ has ranked school $s$ first. Since $s$ has an empty seat under $\beta^{k}(P,\succ,q)$ or has accepted student $j$ with $i\mathrel{\succ_{s}}j$, there are less than $q_{s}$ students who have ranked school $s$ first under $P^{k}$ and have a higher priority than $i$ under $\succ_{s}$. Therefore, $\beta_{i}^{k}(P^{s}_{i},P_{-i},\succ,q)=s$. Since $s\mathrel{P_{i}}\mu(i)$, $i$ is a manipulating student of $\beta^{k}$ at $(P,\succ,q)$.
@@
Part (iii). By part (ii) every manipulating student of $SD^{k}$ is a blocking student of $SD^{k}(P,\succ,q)$. Let $i$ be a blocking student of $\mu=SD^{k}(P,\succ,q)$. Then at $i$’s turn there is no seat left among her top $k$ acceptable schools. She is then left unmatched while she has ranked a school $s$ as acceptable and below the position $k$ which still has a seat available. Let $P^{s}_{i}$ be a preference relation where $i$ ranks $s$ first. Then $SD^{k}_{i}(P^{s}_{i},P_{-i},\succ,q)=s$. Therefore, student $i$ is a manipulating student of $SD^{k}$ at $(P,\succ,q)$. ∎

### v4

@@ 5. Conclusions
We now prove that $\beta\geq\alpha$. We prove the claim that there is no student in $I\setminus\{i,i_{1},\ldots,i_{n+1}\}$ who is not a blocking student of $\nu$ but a blocking student of $\mu$. Let $j\notin I\setminus\{i,i_{1},\ldots,i_{n+1}\}$ and note that $\nu(j)=\mu(j)$. If student $j$ is matched under $\nu$, then she is not a blocking student of $\nu$ and $\mu$. Suppose that $\nu(j)=\emptyset$, and for a contradiction, that she is not a blocking student of $\nu$ but a blocking student of $\mu$ under $(P,\succ,q)$. There is a school $s^{\prime}$ such that $s^{\prime}\mathrel{P_{j}}\mu(j)$ and either (i) school $s^{\prime}$ has an empty seat under $\mu$ or (ii) there is a student $j^{\prime}$ such that $\mu(j^{\prime})=s^{\prime}$ and $j\mathrel{\succ_{s^{\prime}}j^{\prime}}$. Consider (i). As above, school $s^{\prime}$ has an empty seat under $\nu$. Therefore, student $j$ is also a blocking student of $\nu$, contradicting our assumption. Consider (ii). Suppose that school $s^{\prime}$ did not reject any student in any step from the step at which student $i$ applies to school $s$ to the end. Then $\mu(j^{\prime})=\nu(j^{\prime})=s^{\prime}$ and student $j$ is also a blocking student of $\nu$. Suppose that school $s^{\prime}$ has rejected a student $j^{\prime\prime}$ such that $\nu(j^{\prime\prime})=s^{\prime}$. Since $\mu(j^{\prime})=s^{\prime}$, then $j^{\prime}\mathrel{\succ_{s^{\prime}}}j^{\prime\prime}$. Thus $j\mathrel{\succ_{s^{\prime}}}j^{\prime\prime}$ and since $\nu(j^{\prime\prime})=s^{\prime}$, student $j$ is a blocking student of $\nu$, contradicting again our assumption. Therefore there is no student who is simultaneously not a blocking student of $\nu$ but a blocking student of $\mu$ under $(P,\succ,q)$. Thus $\alpha\geq\beta$.
@@
By this claim, $\alpha+y\geq\beta+x$. Therefore, $\nu$ has a weakly larger number of blocking students than $\mu$. Starting from $N=\emptyset$ and successively replacing students’ preferences in any order we conclude that $GS(P^{k-1},\succ,q)$ has weakly more number of blocking students than $GS(P^{k},\succ,q)$.
@@
Finally, we describe a problem where the outcome of $GS^{\ell}$ has more blocking students than the outcome of $GS^{k}$. Let $(P,\succ,q)$ be a problem where each school has one seat, each student has $k$ acceptable schools, and such that students have a common ranking of schools. Then, $GS^{k}(P,\succ,q)=GS(P,\succ,q)$. Thus $GS^{k}(P,\succ,q)$ is stable at $(P,\succ,q)$. Let $s$ be the school that students have ranked at the $k$’th position starting from the top. Since there are more students than schools and $k>\ell$, at least one student is not matched under $GS^{\ell}(P,\succ,q)$ and no student is matched to school $s$ even though every student prefers it to be unmatched. Therefore, there are more blocking students for $GS^{\ell}(P,\succ,q)$ than $GS^{k}(P,\succ,q)$ under $(P,\succ,q)$. ∎
@@
Part (i): Let $i$ be a blocking student of $\mu=\beta(P^{k},\succ,q)$. There is a school $s$ such that the pair $(i,s)$ blocks $\mu$ under $(P,\succ,q)$. Then, $s\mathrel{P_{i}}\mu(i)$ and either (a) school $s$ has an empty seat under $\mu$ or (b) there is a student $j$ such that $\mu(j)=s$ and $i\mathrel{\succ_{s}}j$. We claim that student $i$ did not rank school $s$ first under $P_{i}$. Otherwise, school $s$ has rejected student $i$ at the first step of the Boston algorithm under $(P^{k},\succ,q)$. This is because $k>1$ and the top-ranked schools are considered under $\beta^{k}$. This contradicts the assumption that school $s$ has an empty seat or has accepted student $j$ with $i\mathrel{\succ_{s}}j$. Let $P^{s}_{i}$ be a preference relation where $i$ has ranked school $s$ first. Since $s$ has an empty seat under $\beta^{k}(P,\succ,q)$ or has accepted student $j$ with $i\mathrel{\succ_{s}}j$, there are less than $q_{s}$ students who have ranked school $s$ first under $P^{k}$ and have a higher priority than $i$ under $\succ_{s}$. Therefore, $\beta_{i}^{k}(P^{s}_{i},P_{-i},\succ,q)=s$. Since $s\mathrel{P_{i}}\mu(i)$, $i$ is a manipulating student of $\beta^{k}$ at $(P,\succ,q)$.
@@
Part (iii). By part (ii) every manipulating student of $SD^{k}$ is a blocking student of $SD^{k}(P,\succ,q)$. Let $i$ be a blocking student of $\mu=SD^{k}(P,\succ,q)$. Then at $i$’s turn, there is no seat left among her top $k$ acceptable schools. She is then left unmatched while she has ranked a school $s$ as acceptable and below the position $k$ which still has a seat available. Let $P^{s}_{i}$ be a preference relation where $i$ ranks $s$ first. Then $SD^{k}_{i}(P^{s}_{i},P_{-i},\succ,q)=s$. Therefore, student $i$ is a manipulating student of $SD^{k}$ at $(P,\succ,q)$. ∎